OpenStudy (anonymous):

Simplify the expression below. 6n^2 - 5n^2 + 7n^2

OpenStudy (hoblos):

they are all like terms so just add the coefficients 6n^2 - 5n^2 + 7n^2 = (6-5+7)n^2 = 8n^2